The probability that a family has children is 0.73, the probability that a family has children and owns an iPhone is 0.06, and the probability that a family has children or owns an iPhone is 0.85. Calculate the probability a family owns an iPhone.

Answer by ikleyn(52879)

.
The probability that a family has children is 0.73, the probability that a family has children and
owns an iPhone is 0.06, and the probability that a family has children or owns an iPhone is 0.85.
Calculate the probability a family owns an iPhone.
~~~~~~~~~~


    P(or) = P(C) + P(i) - P(and)


    0.85  = 0.73 + 0.06 - P(and)


    P(and) = 0.73 + 0.06 - 0.85 = 0.79 - 0.85 = -0.06.



The probability can not be negative  ---  HENCE, the problem is posed  INCORRECTLY.


This problem is a  FAKE  or a  TRAP,  or  BOTH.

It describes a case which  NEVER  may happen.